AnimeGhoul8863:

Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation, and determine if the supervisor is correct or not.

AnimeGhoul8863:

@Hero Help preaze e.e

AnimeGhoul8863:

@Shadow

AnimeGhoul8863:

@dude

Hero:

Well, I really just use an equation to solve this one

Hero:

Let \(A\) represent the time it takes person \(A\) to do the job alone and \(B\) represent the time it takes person \(B\) to do the job alone. Then \(t\) represents the time it takes both \(A\) and \(B\) to do the job together as represented by: \(t = \dfrac{AB}{A + B}\)

Hero:

So this this case, we can actually use \(J\) and \(M\) for variables

Hero:

\(t = \dfrac{JM}{J + M}\) where \(J\) represents how fast Jean can fill the pool and \(M\) represents how fast Mark an fill the pool.

AnimeGhoul8863:

okie so it would look like..... \[t=\frac{ (8)(12) }{ 8+12 }\]

AnimeGhoul8863:

like dis???

Hero:

Very nice :)

Hero:

Yes correct. Now try to solve for \(t\) in fraction form. Don't reduce to decimals.

Hero:

Try to write it as a mixed fraction.

AnimeGhoul8863:

okie let me see here

AnimeGhoul8863:

the only thing i can think of is \[4\frac{ 16 }{ 20}\] ?????? this is a guess and the only thing i can think of

Hero:

Reduce the fraction

Hero:

It's better to write it as \(4 + \dfrac{16}{20}\)

Hero:

Writing it the other way implies multiplying 4 times 16

AnimeGhoul8863:

ok so \[4+\frac{ 16 }{ 20 }\] \[\frac{ 16 }{ 20 }= \frac{ 4 }{ 5 }\] so it would be \[4+\frac{ 4 }{ 5 }\] convert it to fraction \[\frac{ 4*5 }{ 5}+\frac{ 4 }{ 5 }\]

AnimeGhoul8863:

am i correct so far?

Hero:

\(4 + \dfrac{4}{5}\) is good enough

AnimeGhoul8863:

oh okie e.e

Hero:

Now you have to convert the 4/5 to minutes using a proportion

Hero:

You have 4 hours and then 4/5 of an hour but the 4/5ths is meaningless because it doesn't represent a full hour so you have to convert it to minutes using the following proportion: \(\dfrac{4}{5} = \dfrac{x}{60}\)

AnimeGhoul8863:

so it would take a third of the time if they did it together

Hero:

\(4 + \dfrac{4}{5}\) means it will take four hours plus a fraction of an hour. We have to interpret what that fraction of an hour is in minutes. So solve the proportion above.

AnimeGhoul8863:

OOHHH i think i got it \[\frac{ 4 }{ 5 }=\frac{ 48 }{ 60}\] i could be wrong but i think this is the answer im looking for

Hero:

Looks right to me. So basically it takes 4 hours and 48 minutes for them to do the same job working together. These problems are a lesson in the power of working together to accomplish tasks. You could do it alone but it would take much longer. If I had to do a job and I was told I had 12 hours to do it but my good ole buddy pal is with me and we could easily do it faster if we worked together, I'd rather take the fastest route and save 7 hours.

AnimeGhoul8863:

so all together it would look like \[t=\frac{ JM }{ J+M }\] \[t=\frac{ (8) (12)}{ 8+12 }\] Write in a mixed fraction \[4+\frac{ 16 }{ 20 }\] \[reduce \frac{ 16 }{ 20 }=\frac{ 4 }{ 5 }\] \[4+\frac{ 4 }{ 5 }\] \[\frac{ 4 }{ 5}=\frac{ x }{ 60 }\] \[x=48\] \[\frac{ 4 }{ 5}=\frac{ 48 }{ 60 }\] Meaning the time it takes to fill the pool working together would take 4 hours and 48 minutes

Hero:

Well, you'd still have to put explanations in full sentences at each step.

Hero:

Which should not be difficult for you

AnimeGhoul8863:

right but over all this is the answer and how it should look

Hero:

Yes

AnimeGhoul8863:

thank you! Is there anyway you could help with one more thats like this one

Hero:

Only thing is, your teacher might ask you where you got the equation. Just say you got it from Hero

AnimeGhoul8863:

x'D you got it

Hero:

She might ask how did you derive it though. Which would be even worse

AnimeGhoul8863:

i dont think so my teacher just see's the right answer and there like "pass" he doesnt ask much questions like that

Hero:

Great job on this btw. I can tell you're eager to learn and teachable